OpenStudy (anonymous):

Evaluate the Indefinite Integral:

OpenStudy (anonymous):

\[\int\limits_{?}^{?} dx \div \sqrt{x} (1-\sqrt{x})^{3}\]

OpenStudy (anonymous):

i got 6 mins as my answer

OpenStudy (anonymous):

make the u sub of u=(1-sqrt(x)) du =-1/(2*sqrt(x)) and you end up with\[-2\int\limits_{?}^{?} u^{-3} du \] and the rest should be simple

OpenStudy (anonymous):

That's what I got as my first few steps - did you get 1 / 1-\[1 / (1- \sqrt{x})^{2}\]

OpenStudy (anonymous):

yea thats right

OpenStudy (anonymous):

Fantastic! Thanks!!

OpenStudy (anonymous):

Did you get the same exponent?

OpenStudy (anonymous):

yea exactly the same